The Manufacture of television tubes knows from the past experience that the average life of tube is 2000hrs with a s.d. of 200 hrs. A sample of 100 tubes has average life of 1950 hrs. Test at the 0.01 level of significance to see if this sample came from a normal population of mean 2000 hrs.


1
Expert's answer
2021-12-06T17:40:28-0500

The following null and alternative hypotheses need to be tested:

"H_0: \mu=2000"

"H_1:\mu\not=2000"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a two-tailed test is "z_c=2.5758."

The rejection region for this two-tailed test is "R = \{z: |z| > 2.5758\}."

The z-statistic is computed as follows:



"z=\dfrac{\bar{X}-\mu}{\sigma/\sqrt{n}}=\dfrac{1950-2000}{200/\sqrt{100}}=-2.5"

Since it is observed that "|z| = 2.5 \le 2.5758=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach: The p-value is "p=2P(Z<-2.5)\approx0.012419," and since "p=0.012419>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is different than "2000," at the "\alpha = 0.01" significance level.

Therefore, there is enough evidence to claim that this sample came from a normal population of mean 2000 hrs, at the "\alpha = 0.01" significance level.